OpenStudy (anonymous):

What is the slope and y-intercept of the graph of -x + 2y = 6?

OpenStudy (anonymous):

First, put it into slope-intercept form which is y=mx+b. So, -x+2y=6 2y=x+6 /2 /2 y=1/2x+3 "m" represents slope while "b" represents y-intercept. m=1/2 b=3

OpenStudy (anonymous):

Thanks, sorry didn't respond soon.